### v2

# Reaching Curzon-Ahlborn limit in linear response and Whitney limit in nonlinear response in edge mode quantum thermoelectrics and refrigeration

Abstract
Quantum heat engines and quantum refrigerators are proposed in three-terminal quantum Hall (QH) and quantum spin Hall (QSH) setups with a voltage-temperature probe in both the linear and nonlinear transport regimes. In the linear response regime, we find that efficiency at maximum power approaches the Curzon-Ahlborn limit in both QH and QSH setups. Similarly, in nonlinear response, we find that efficiency at maximum power reaches the Whitney bounds. This is for the first time, we see that in the same setup and using quantum point contacts, the thermoelectric efficiency limits in linear and nonlinear response being achieved.
preprint: APS/123-QED
Introduction: The quantum Hall resistance [1] in two-dimensional electron gas in the presence of a normal uniform magnetic field at low temperature arises as a consequence of chiral edge modes [2], which is famously known as quantum Hall effect. These edge modes are robust against disorder and are immune to backscattering [3] being topologically protected. Complimentary to chiral edge modes, helical edge modes are seen in quantum spin Hall experiments with Mercury Telluride/Cadmium Telluride heterostructures [4, 5, 6], which are also topologically protected. Recently in QH [7, 8, 9, 10, 11] and QSH [12, 13] setups, thermoelectric properties have been investigated using Landaeur-Buttiker formalism [14, 15]. Quantum transport can be described in two regimes. First is linear, wherein the applied voltage and temperature biases are quite small, and thus leads to a linear relation between current and the temperature or voltage biases applied [16]. The second is nonlinear, wherein the applied voltage and temperature biases are not small. In the linear response regime, the efficiency at maximum power ($\eta|_{P_{max}}$) and the maximum efficiency ($\eta_{max}$) are closely related to the figure of merit $Z\theta\hskip 2.84544pt\left(=\frac{GS^{2}\theta}{K}\right)$, where $G$ = conductance, $S$ = Seebeck coefficient, $K$ = thermal conductance and $\theta$ is the reference temperature of the setup. In this regime, there exists a bound to $\eta|_{P_{max}}$, which is given by Curzon-Ahlborn (CA) [17] limit with the efficiency ($\eta|_{P_{max}}$)= $\eta_{c}/2$, where $\eta_{c}=1-\theta_{2}/\theta_{1}$ is the Carnot efficiency with $\theta_{1}$ and $\theta_{2}$ being the temperatures of the hot and cold reservoirs respectively [16]. To achieve this, it is necessary to have a large $Z\theta$. This can be accomplished by ensuring a high Seebeck coefficient ($S$) and low thermal conductance ($K$). Similarly, in the non-linear transport regime, there also exists an upper bound to $\eta|_{P_{max}}$, the Whitney limit [18, 19], i.e., efficiency at maximum power output ($\eta|_{P_{max}^{Wh}}$) = $\eta|_{P_{max}^{Wh}}$ = $\eta_{c}/(1+0.936(1+\theta_{2}/\theta_{1}))$, $N$ being the number of edge modes. This is the nonlinear equivalent of the CA efficiency in linear response and has been derived utilizing BoxCar-type transmission. The study of two-terminal setups in nonlinear response was extended to a three-terminal setup with a voltage probe, and it was shown that the Whitney limits are universal [20]. Similarly, in [21], it has been proved that a generic mesoscopic thermoelectric heat engine can be highly efficient only when the transmission is Box-car type. Boxcar-type scattering can be implemented via a chain of quantum dots [18, 19, 20]. However, electron-electron interaction between quantum dots, which can lead to coulomb blockade effects, is neglected in the calculation in Refs. [18, 19, 20]. This assumption is quite extreme, as in real systems, having a chain of quantum dots without the coulomb blockade effect is almost impossible to implement [16]. In contrast, this work considers quantum point contact (QPC) type scattering, which is easier to implement experimentally [22, 23, 24, 25] than the boxcar-type transmission. In this work, we reach the Whitney bound in the nonlinear transport regime using QPC-type scattering. Previously, in Refs. [18, 19, 20] a QH setup was only considered; however, in this letter, we consider both QH and QSH setups, albeit with the easier-to-implement QPC type contacts for both linear as well as non-linear response regimes.
Figure 1: (a) 3T QH sample, (b) 3T QSH sample with a VT probe and two QPC-type constrictions connected to an external gate voltage ($V_{g}$) in either sample. Purple lines depict electron edge modes in QH samples, while those in QSH samples are shown by purple and red lines for spin-up and spin-down electrons.
This letter examines the thermoelectric performance of three-terminal QH and QSH setups using a voltage-temperature (VT) probe in both linear and nonlinear regimes (see Fig. 1). The VT probe incorporates inelastic scattering processes from electron-electron or electron-phonon interactions, ensuring no net charge ($I_{3}=0$) and heat current ($J_{3}=0$) through terminal 3. This is more accurate than a voltage probe, which insists on a vanishing net charge current [12, 7]. A VT probe  [26] has been used to compare the Lorentz ratio scaling with the Luttinger-liquid model [27]. This letter proposes three-terminal QH and QSH setups with the third terminal as a VT probe, functioning as a highly efficient quantum heat engine (QHE) or quantum refrigerator (QR) in both linear and nonlinear response regimes.
This letter is divided into two parts. In the first part, we investigate the thermoelectric transport of the QH and QSH setups within the linear response regime and show that efficiency at maximum power approaches Curzon-Ahlborn for both QH and QSH setups. In the second part, we focus on non-linear transport, wherein we see that the efficiency at maximum power approaches the Whitney limit for both QH and QSH setups. The technical details of the calculation are presented in the supplementary material (SM).
Thermoelectricity in QH and QSH setups: We start by discussing the general theory for thermoelectric transport in QH as well as QSH setups within the Landaeur-Buttiker scattering theory in both linear and non-linear regimes [16, 14]. The net charge and heat currents in a multiterminal QH sample, see Sec. II of SM, are given as,
[…]
where the transmission probability for an electron to scatter from terminal $\beta$ to terminal $\alpha$ is $T_{\alpha\beta}$. Here, $f_{\alpha(\beta)}=\left(1+e^{(E-\mu_{\alpha(\beta)})/k_{B}\theta_{\alpha(\beta)}}\right)^{-1}$, $\mu_{\alpha(\beta)}$ and $\theta_{\alpha(\beta)}$ are the Fermi function, chemical potential and temperature of terminal $\alpha(\beta)$ respectively.
Similarly, for QSH setup, with helical edge modes, the currents carry an extra spin degree of freedom and are written as $I_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{\alpha}^{\sigma}$ and $J_{\alpha}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{\alpha}^{\sigma}$, where [12],
$\begin{split}I_{\alpha}^{\sigma}&=\frac{e}{h}\int_{-\infty}^{\infty}dE\sum_{\beta}T_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\\ J_{\alpha}^{\sigma}&=\frac{1}{h}\int_{-\infty}^{\infty}dE(E-\mu_{\alpha})\sum_{\beta}T_{\alpha\beta}^{\sigma}(f_{\alpha}(E)-f_{\beta}(E))],\end{split}$ (2)
with the transmission probability for an electron to transmit from terminal $\beta$ with spin $\rho$ to terminal $\alpha$ with spin $\sigma$ being $T_{\alpha\beta}^{\sigma}=\sum_{\rho\in\{\uparrow,\downarrow\}}T_{\alpha\beta}^{\sigma\rho}$. In this three-terminal setup (both QH and QSH) as shown in Fig. 1, $\theta_{1}=\theta+\tau_{1},\theta_{2}=\theta+\tau_{2}$ and $\theta_{3}=\theta+\tau_{3}$, where $\theta$ is the reference temperature and $\tau_{1},\tau_{2}$ and $\tau_{3}$ are the temperature biases applied in terminals 1, 2 and 3 respectively. For our setups, we apply temperature bias only at terminal 1, i.e., $\tau_{1}=\tau$ and $\tau_{2}=0$. We study the thermoelectric properties for a QHE, considering terminal 3 to be a VT probe. We also assume both terminals 1 and 2 to be current probes with $V_{1}=-V$ and $V_{2}=0$. Here, we have considered only one type of energy-dependent transmission via constrictions, which is known as the quantum point contact (QPC) type [28], defined via $T_{\text{QPC}}=[1+\text{exp}(-2\pi(E-E_{l})/\hbar\omega]^{-1}$, where $l$ denotes the constriction index, i.e., $l\in\{1,2\}$. Below a certain threshold energy level $E_{l}$, $T_{QPC}$ vanishes, while above $E_{l}$, it is transparent, while $\hbar\omega$ indicates the width of the transmission resonance, see also Sec. I of SM. Following Refs. [29, 30], a capacitive external gate voltage $V_{g}$ is connected to the QPC constrictions which, in turn, controls the threshold energy $E_{l}$ of QPC, through which we can explore both linear and nonlinear transport regimes for our setups in Fig. 1(a) and (b).
Linear transport in QH and QSH setups: In the linear response regime for the QH case, i.e., when $\mu_{\alpha}(=eV_{\alpha})\ll k_{B}\theta$ and ($\theta_{\alpha}-\theta)\ll\theta$, one can see that the current (both charge and heat) varies linearly with the voltage and temperature biases applied across the sample [16]. Since, terminal 3 is a VT probe i.e., $I_{3}=J_{3}=0$, the charge and heat currents in terminal 1 can be written as,
$\begin{pmatrix}I_{1}\\ J_{1}\end{pmatrix}=\begin{pmatrix}L_{eV}&L_{e\theta}\\ L_{hV}&L_{h\theta}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (3)
where the Onsager matrix elements are $L_{eV},L_{e\theta},L_{hV},L_{h\theta}$, see Eq. (31) in SM. These Onsager matrix elements are used to find the transport coefficients such as conductance ($G$), Seebeck coefficient ($S$), Peltier coefficient ($\Pi$), and the thermal conductance ($K$) and are given as,
$G$ $=L_{eV},\quad S=\frac{L_{e\theta}}{L_{eV}},\quad\Pi=\frac{L_{hV}}{L_{eV}},\quad K=L_{h\theta}-\frac{L_{hV}L_{e\theta}}{L_{eV}}.$ (4)
Using a similar procedure, one can also derive the Onsager matrix elements in the QSH case for charge current $I_{\alpha}^{\sigma}$ and heat current $J_{\alpha}^{\sigma}$ by imposing the VT probe condition ($I_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}I_{3}^{\sigma}=0$ and $J_{3}=\sum_{\sigma\in\{\uparrow,\downarrow\}}J_{3}^{\sigma}=0$) and they can be written in the same form as Eq. (3) (see, Sec. III.1.1 in SM) as
$\begin{pmatrix}I_{1}^{\sigma}\\ J_{1}^{\sigma}\end{pmatrix}=\begin{pmatrix}L_{eV}^{\sigma}&L_{e\theta}^{\sigma}\\ L_{hV}^{\sigma}&L_{h\theta}^{\sigma}\end{pmatrix}\begin{pmatrix}-V\\ \tau\end{pmatrix}$ (5)
where the spin polarized Onsager matrix elements $L_{eV}^{\sigma}$, $L_{e\theta}^{\sigma}$, $L_{hV}^{\sigma}$ and $L_{h\theta}^{\sigma}$ are derived in Eq. (65) of SM. Using Eq. (5), the spin polarized transport coefficients such as Conductance ($G^{\sigma}$), Seebeck Coefficient ($S^{\sigma}$), Peltier Coefficient ($\Pi^{\sigma}$), and thermal conductance ($K^{\sigma}$) are defined as,
$G^{\sigma}$ $=L_{eV}^{\sigma},\quad S^{\sigma}=\frac{L_{e\theta}^{\sigma}}{L_{eV}^{\sigma}},\quad\Pi^{\sigma}=\frac{L_{hV}^{\sigma}}{L_{eV}^{\sigma}},\quad K^{\sigma}=L_{h\theta}^{\sigma}-\frac{L_{hV}^{\sigma}L_{e\theta}^{\sigma}}{L_{eV}^{\sigma}}.$ (6)
The output power in terminal 1, $P=I_{1}V$ is maximum at a voltage bias $V=L_{e\theta}\tau/4L_{eV}$ and is given as $P_{max}=L_{e\theta}^{2}\tau^{2}/4L_{eV}$, see Sec. II.1.2 for QH and III A 2 for QSH in SM. The efficiency at maximum power, i.e., $\eta|_{P_{max}}=\frac{P_{max}}{J}=\frac{\eta_{c}}{2}\frac{Z\theta}{Z\theta+2}$ (see, Eq. (34) for QH and Eq. (75) for QSH in SM), $J=J_{1}$ is the heat current from the hotter terminal 1. $Z\theta=\frac{GS^{2}\theta}{K}$ is the figure of merit. Similarly, maximum efficiency, i.e., $\eta_{max}=\eta_{c}\frac{\sqrt{Z\theta+1}-1}{\sqrt{Z\theta+1}+1}$ can be derived (see Eq. (38) for QH and Eq. (79) for QSH in SM).
In Fig. 2(a, b), we parametrically plot the efficiency ($\eta/\eta_{c}$) and the power ($P/P_{max}$) in linear response regime for both QH and QSH QHE, calculated using Eqs. (3-6) and observe that it achieves the Curzon-Ahlborn limit at $V=0.057k_{B}\theta/e$ and further increasing $V$, power decreases and efficiency increases (see blue (red) dashed curve for $V=0.081k_{B}\theta/e$ and blue (red) dotted curve for $V=0.096k_{B}\theta/e$) . We considered threshold energies of QPC 1 and QPC 2 to be $E_{1}^{{}^{\prime}}=E_{1}+eV_{g}$ and $E_{2}^{{}^{\prime}}=E_{2}+eV_{g}$ respectively, with $E_{1}=E_{2}$. If $E_{1}\neq E_{2}$; the conclusions don’t change, see, Secs. II.1.4 (II.2.4) for QH and III A 3 (III B 4) for QSH for the linear (non-linear) transport in SM.
In Ref. [7], it has been reported that a QH setup cannot work as a QR due to time-reversal symmetry breaking (TRS). Conversely, the same setup with QSH edge modes can act as both QHE and QR because TRS is preserved [12]. In the QH setup as in [7], the asymmetric parameter (AP), i.e., the Seebeck to Peltier coefficient ratio, is zero or infinity, affecting its refrigeration capability [31, 12]. In the QSH setup, AP is always finite since TRS is preserved [12]. However, this is not true for all QH setups. The QH setup in Fig. 1(a) and the QSH setups in Fig. 1(b) can act as both QHE and QR because their AP is finite (see Sec. II.1.4 of SM).
For QR, we can calculate the cooling power at the maximum coefficient of performance ($\textbf{J}|_{\eta^{r}_{max}}$) and the maximum coefficient of performance ($\eta_{max}^{r}$) for both QH and QSH setups. The coefficient of performance (COP = $\textbf{J}/P$) is generally defined as the ratio of heat extracted from the cooler terminal ($\textbf{J}=-J_{2}$) to the power absorbed ($P$) by the setup. The maximum COP ($\eta^{r}_{max}$) is achieved using the condition $\frac{d(COP)}{dV}=0$ and is given as $\eta^{r}_{max}=\eta^{r}_{c}\frac{\sqrt{Z\theta+1}-1}{\sqrt{Z\theta+1}+1}$, where $\eta^{r}_{c}$ is the Carnot coefficient of performance. As shown in Fig. 2(c), $\eta^{r}_{max}$ reaches $0.97\eta_{c}^{r}$ at $\textbf{J}|_{\eta^{r}_{max}}$ at $V=0.115k_{B}\theta/e$. Similar to the QH setup, we parametrically plot the cooling power J vs coefficient of performance $\eta^{r}$ in Fig. 2(d) for QSH setup.
As shown in Sec. II.1.2, for QH and Sec. III.1.2 for QSH setups in the SM, both $\eta_{max}$ and $\eta|_{P_{max}}$ rely on $Z\theta$ for a QHE. Similarly, for QR, $\eta^{r}_{max}$ depend on $Z\theta$. To achieve Curzon-Ahlborn efficiency ($\frac{\eta_{c}}{2}$), $Z\theta$ must significantly exceed 1. The key lies in having a large Seebeck coefficient ($S$) and a small thermal conductance ($K$), both occurring in the same configuration. In both of our QH and QSH setups, we find that the Seebeck coefficient is notably large, while the thermal conductivity is remarkably small..
(a) QH heat engine (linear)
(b) QSH heat engine (linear)
(c) QH refrigerator (linear)
(d) QSH refrigerator (linear)
Figure 2: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}}$ (a) QH, (b) QSH QHE. the blue (red) dot-dashed curve is for $V=0.057k_{B}\theta/e$, the blue (red) dashed curve for $V=0.081k_{B}\theta/e$ and the blue (red) dotted curve for $V=0.096k_{B}\theta/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot of $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}|_{\eta^{r}_{max}}}$ for (c) QH and (d) QSH QR, for QH (QSH) QR, blue (red) line is for $V=0.115k_{B}\theta/e$. Parameters taken are $\omega=0.1k_{B}\theta/\hbar$, $\theta_{1}=1.1K,\theta_{2}=1.0K$, $\mu=0$, $E_{1}=E_{2}=k_{B}\theta$.
(a) QH heat engine (nonlinear)
(b) QSH heat engine (nonlinear)
(c) QSH refrigerator (nonlinear)
(d) QSH refrigerator (nonlinear)
Figure 3: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}^{Wh}}$ for (a) QH and (b) QSH, QHE. For QH (QSH) QHE, the blue (red) dot-dashed curve is for $V=1.14k_{B}\theta/e$, the blue (red) dashed curve for $V=1.77k_{B}\theta/e$ and the blue (red) dotted curve for $V=3.7k_{B}\theta/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot for $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}^{Wh}_{max}}$ for (c) QH and (d) QSH, QR. For QH (QSH) QR, blue (red) line is for $V=20k_{B}\theta/e$. Parameters taken are $\omega=\frac{0.1k_{B}\theta}{h}$, $\theta_{1}=2K,\theta_{2}=1K$, $\mu=0$, $E_{1}=E_{2}=k_{B}\theta$.
Non-linear Transport in QH and QSH Setups: There has been ongoing work in this direction on nonlinear thermoelectric transport [18, 19, 32, 33, 34] in both QH and QSH setups. In a two-terminal QH setup, significant deviations from the Onsager-Casimir relation have been observed [32] with a similar investigation done in QSH setup [33]. In a three-terminal QH setup [34] with a voltage probe, efficient operation as a QR in the weakly non-linear regime has been demonstrated. Additionally, a normal metal Aharonov-Bohm heat engine exhibits enhanced thermoelectric performance in the nonlinear regime [29], surpassing its linear regime counterpart [30].
In the linear response regime, electron transport depends solely on the electron’s kinetic energy. However, in the nonlinear regime, it relies on both the kinetic energy and the interaction potential ($U^{\prime}$) of the sample where the electrons flow. It implies that whereas in the linear regime, the QPC energy level $E_{l}$ shifts via the gate voltage: $E_{l}\longrightarrow E_{l}+eV_{g}$, in the nonlinear regime, the interaction potential $U^{\prime}$ shifts the QPC energy level $E_{l}$ as $E_{l}\longrightarrow E_{l}+eU^{\prime}$, with $U^{\prime}=U+u_{g}V_{g}$, where $u_{g}$ is the characteristic potential determining the response of $U^{\prime}$ to the gate voltage $V_{g}$. For the form of $U$, see Eq. (54) of SM. Introducing a gate voltage helps nullify the interaction potential ($U$) in the nonlinear regime, as explained in Sec. II.2.2 of SM and Ref. [29]. From discussion in Sec. II.2.2 of SM , the gate voltage must be of the form $V_{g}=(-U+V_{g}^{\prime})/u_{g}$, where the first term inside the parentheses nullifies the interaction potential $U$, while the second term raises the energy level $E_{l}$: $E_{l}\longrightarrow E_{l}+eV_{g}^{\prime}$ for the nonlinear response regime. We have given an estimation for the interaction potentials in two-terminal QH and three-terminal QH setups with a voltage-temperature probe in SM (see, Tables I and II of SM) for different applied biases. Similarly, we provide an estimate of the interaction potential for two-terminal QSH and three-terminal QSH setups with voltage-temperature probe in SM (Tables III and IV of SM).
For nonlinear transport, we start with Eq. (1) for QH and (2) for QSH, respectively. Unlike linear response, solving this analytically is challenging, but numerical solutions are possible. In the Mathematica code [35], for QHE, we numerically solve for $V_{3}$ and $\tau_{3}$ at a given $V$ and $\tau$, satisfying the VT probe condition ($I_{3}=J_{3}=0$), then use these values to determine the current $I_{1}$ and $J_{1}$ (via Eqs. (1) and (2) for QH and QSH respectively), and finally compute power $P=VI_{1}$ and efficiency $\eta=\frac{P}{J_{1}}$. Similarly, for QR, we determine the cooling power $\textbf{J}=-J_{2}$ and the coefficient of performance $\eta^{r}=\frac{\textbf{J}}{P}$, where $P$ is the power absorbed by the setup.
In QH setups, as shown in Refs. [18, 19], there exists a bound to the maximum power ($P_{max}=P_{max}^{Wh}$), given by $0.0642\pi^{2}Nk_{B}^{2}(\theta_{1}-\theta_{2})^{2}/h$. Similarly, the bound to the efficiency at maximum power ($\eta|_{P_{max}}$) is found at $P_{max}^{Wh}$, yielding $\eta|{P_{max}^{Wh}}=\eta_{c}/(1+0.936(1+\theta_{2}/\theta_{1}))$, where $\theta_{1}$ and $\theta_{2}$ are the temperatures of the hot and cold reservoirs, and $N$ is the number of edge modes. In our case, with $\theta_{1}=2K$, $\theta_{1}=1K$, and $N=1$, we have $P_{max}^{Wh}=0.632k_{B}^{2}/h$ and $\eta|_{P_{max}^{Wh}}=0.41\eta_{c}$. For power output $P\ll P_{max}^{Wh}$, the bound to $\eta_{max}$ is given by $\eta|_{max}^{Wh}=\eta_{c}\left(1-0.478\sqrt{\theta_{2}P/\theta_{1}P_{max}^{Wh}}\right)$. These quantities were derived for a two-terminal QH setup with a constriction and boxcar-type transmission (see, Sec. II.2.1 of SM and also Refs. [18, 19]). We extend the same approach to a two-terminal QSH setup and prove that the bounds to $P_{max}$, $\eta|_{P_{max}}$, and $\eta_{max}$ are exactly similar to the QH case (see, Sec. II.2.1 of SM).
[…]
Table 1: Thermoelectric performance of the QH and QSH setup in the linear and nonlinear response regime. $P$ is in units of $P_{max}$ for linear, and $P_{max}^{Wh}$ for nonlinear, J is in units of $\textbf{J}|_{\eta^{r}_{max}}$ for linear and $J_{max}^{Wh}$ for nonlinear.
QHE QR
Regime Setup $P$ $\frac{\eta|_{P_{max}}}{\eta_{c}}$ $\frac{\eta_{max}}{\eta_{c}}$ J $\frac{\eta^{r}_{max}}{\eta^{r}_{c}}$
Linear QH
1
[Fig. 2(a)]
0.50
[Fig. 2(a)]
0.97
[Fig. 2(a)]
1
[Fig. 2(c)]
0.97
[Fig. 2(c)]
QSH
1
[Fig. 2(b)]
0.50
[Fig. 2(b)]
0.97
[Fig. 2(b)]
1
[Fig. 2(d)]
0.97
[Fig. 2(d)]
Nonlinear QH
1
[Fig. 3(a)]
0.41
[Fig. 3(a)]
0.93
[Fig. 3(a)]
1
[Fig. 3(c)]
0.86
[Fig. 3(c)]
QSH
0.76
[Fig. 3(b)]
0.41
[Fig. 3(b)]
0.93
[Fig. 3(b)]
0.75
[Fig. 3(d)]
0.86
[Fig. 3(d)]
Adjusting the gate voltage to nullify the interaction potential shifts the energy level of the QPC-like constriction. This is explained in Sec. II.2.2 of SM, initially demonstrated in a two-terminal nonlinear QH setup and extended to a three-terminal configuration (Sec. II.2.4 of SM). A similar approach is applied to a two-terminal QSH setup in Sec. III.2.2 of SM, later extended to a three-terminal configuration (Sec. III.2.4 of SM). We examine $P_{max}$, $\eta|_{P_{max}}$ and $\eta_{max}$ numerically. In Fig. 3(a), for the QH setup, maximum power approaches the Whitney limit at a specific voltage bias $V=1.14k_{B}\theta/e$ (blue dot-dashed curve). With increasing voltage bias $V$, power decreases and efficiency increases (see blue dashed line for $V=1.77k_{B}\theta/e$ and blue dotted for $V=3.7k_{B}\theta/e$ in Fig. 3(a)). The QHE is more efficient as $\eta|_{P_{max}}$ and $\eta_{max}$ reach $0.41\eta_{c}$ and $0.93\eta_{c}$ respectively. Similarly, in Fig. 3(c), $J_{max}$ reaches $J_{max}^{Wh}$, while the $\eta_{max}^{r}$ is approximately 0.86 $\eta_{c}^{r}$. Details for generating the plots in Fig. 3(a) and (c) are provided in Sec. II.2.4 of SM.
[…]
Experimental Realization and Conclusion: In three-terminal QH and QSH setups with QPC-like constrictions, we achieve optimal thermoelectric performance approaching efficiency benchmarks for QHE and coefficient of performance for QR, both in linear (Fig. 2) and nonlinear (Fig. 3) response regimes. For nonlinear response, maximum power generated approaches Whitney bound $P_{max}^{Wh}$ (see Fig. 3(a), Table 1) and maximum cooling power approaches $J_{max}^{Wh}$ (see Fig. 3(c), Table 1) for QH setup. In QSH, maximum power as QHE decreases from $P_{max}^{Wh}$, yet efficiency at maximum power remains $0.41\eta_{c}$ (Curzon-Ahlborn efficiency equivalent) (see Fig. 3(b), Table 1). Similarly, as QR, maximum cooling power is below $J_{max}^{Wh}$ (Fig. 3(d), Table 1). This power suppression (both $P_{max}^{Wh}$ and $J_{max}^{Wh}$) arises as a result of helical edge mode transport with a VT probe as there is a finite spin current in three-terminal QSH setup, which is absent in chiral case. Designing a VT probe experimentally is slightly challenging due to potential heat flow between reservoir and surroundings, altering its temperature [16]. We employ QPC-type tunneling, utilized in prior works [22, 23, 24, 25] to study electron transport in 2DEGs, and it’s easier to implement experimentally and is controlled electrostatically with split gates [36, 37, 38].
In the linear response regime, an optimal quantum thermoelectric material is characterized by a figure of merit ($Z\theta$) significantly greater than 1. Theoretical proposals report $Z\theta$ values of 4 in ferromagnet-superconductor junctions (Ref. [39]) and up to 40-50 in other studies (Refs. [40], [41]). Materials with $Z\theta\gg 1$ are highly desirable for optimal thermoelectric performance as both quantum heat engines and refrigerators. To summarize, the main take home messages of this letter are: 1. Use of QPC instead of boxcar type transmission [38, 24, 22, 25], which is much easier to implement, 2. Investigating thermoelectric transport in the non-linear response regime for QSH setups, which has never been attempted, 3. Investigation of QH and QSH setups as QHE and QR with a voltage-temperature probe in the non-linear response regime, which has again never been explored before and finally 4. An estimate of interaction potential and the applied gate voltage to nullify it in 2T QSH setup (see, Sec. III.2.2 of SM ) and 3T QH (see, Sec. II.2.4 of SM ) and QSH setups (see, Sec. III.2.4 of SM ) with VTP for the first time.
Quantum heat engines and quantum refrigerators in our work can significantly impact energy conversion and cryogenic cooling. Cooling cryogenic devices to millikelvin temperatures is vital for studying quantum coherence and phase transitions in nano-sized materials (Refs. [16, 42, 43]). Traditionally, cryogenics focused on cooling phonons in solids, but quantum refrigeration now allows more effective electron cooling (Ref. [16]). Our work achieves maximum cooling power, reaching the Whitney bound, implying cooling to the lowest possible temperatures.
In conclusion, this letter proposes achieving the Curzon-Ahlborn (CA) limit in linear and the corresponding Whitney bound in the nonlinear response regimes through a setup applicable to three-terminal QH and QSH systems, extendable to multiterminal setups with multiple VT probes. Furthermore, this approach can be applied to various edge mode systems, including trivial helical [44, 45, 46] and anomalous modes [47, 48, 49].